OpenStudy (anonymous):

What is a solution of x² + 6x = –5? x = –6 x = –1 x = 1 x = 6

OpenStudy (rulnick):

(x+5)(x+1) = 0 x=-5 and x=-1

OpenStudy (rulnick):

... so x=-1.

OpenStudy (anonymous):

x^2 +6x +5 = 0 (x+1)(x+5) = 0 x = -1 x = -5

OpenStudy (anonymous):

they are both right you have to set it to zero then factor it out